OpenStudy (anonymous):

y(x)=f(x+g(x)) Find y'(0) if: f'(1)=1/3 g(0)=1 g'(0)=1/3 can anybody help with this please, I am nearly tearing my hair out, thanks

OpenStudy (jtvatsim):

Well, don't go bald yet. :)

OpenStudy (jtvatsim):

Your first step should probably me to use the chain rule. Have you tried that?

OpenStudy (anonymous):

so f'(x+g(x)).(1+g'(x))?

OpenStudy (jtvatsim):

good!

OpenStudy (jtvatsim):

Now, using the given information what do you get?

OpenStudy (anonymous):

haha give me a moment just thinking

OpenStudy (jtvatsim):

no worries, take your time. :)

OpenStudy (anonymous):

f'((1))+g(0)).(1+g'(0)) =(-1/3+1).(1+1/3)?

OpenStudy (jtvatsim):

close... you should plug in a 0 for all x's (you have a 1 that appears in the first part). We are evaluating y'(0) so x = 0.

OpenStudy (jtvatsim):

\[y'(0) = f'(0 + g(0)) \cdot (1 + g'(0))\]

OpenStudy (jtvatsim):

Does that make sense?

OpenStudy (anonymous):

ah yes ok :)

OpenStudy (jtvatsim):

cool, you are close to slaying this dragon. :)

OpenStudy (anonymous):

so then: f'(0+1).(1+1/3) =f'(1).(1+1/3) =-1/3.(1+1/3)?

OpenStudy (jtvatsim):

In your OP you said that f'(1) = 1/3 positive, but if it should have been negative then yes. :)

OpenStudy (anonymous):

yes sorry it should have been -1/3

OpenStudy (jtvatsim):

Then, you are correct!

OpenStudy (anonymous):

thank you so very very much :)

OpenStudy (jtvatsim):

no worries, and now you can keep your hair. :)

OpenStudy (anonymous):

haha yes :)

OpenStudy (jtvatsim):

take care! nice work!

OpenStudy (anonymous):

awww thanks :), you take care too :)